. Arithmetic mean of 7 numbers is equal to 2.9, and arithmetic mean of other 8 numbers is equal to 2.5. Find the arithmetic mean of all 15 numbers
Answer #1

Mean \bar{x} = \sum X / n

\sum X = \bar{x} * n

For first 7 numbers

\sum X = 2.9 * 7 = 20.3

For other 8 numbers

\sum X = 2.5 * 8 = 20

Therefore, sum of the values of all 15 numbers = 20.3 + 20 = 40.3

Arithmetic mean of 15 numbers = 40.3 / 15

= 2.6867